# C-DOG: Training-Free Multi-View Multi-Object Association in Dense Scenes Without Visual Feature via Connected $\delta$-Overlap Graphs thanks: Code available at GitHub: https://github.com/Yunghong/C-DOG

### v1

@@ Abstract
Multi-view multi-object association is a fundamental step in 3D reconstruction pipelines, enabling consistent grouping of object instances across multiple camera views. Existing methods often rely on appearance features or geometric constraints such as epipolar consistency. However, these approaches can fail when objects are visually indistinguishable or observations are corrupted by noise. We propose C-DOG, a training-free framework that serves as an intermediate module bridging object detection (or pose estimation) and 3D reconstruction, without relying on visual features. It combines connected $\delta$-overlap graph modeling with epipolar geometry to robustly associate detections across views. Each 2D observation is represented as a graph node, with edges weighted by epipolar consistency. A $\delta$-neighbor-overlap clustering step identifies strongly consistent groups while tolerating noise and partial connectivity. To further improve robustness, we incorporate Interquartile Range (IQR)-based filtering and a 3D back-projection error criterion to eliminate inconsistent observations. Extensive experiments on synthetic benchmarks demonstrate that C-DOG outperforms geometry-based baselines and remains robust under challenging conditions, including high object density, without visual features, and limited camera overlap, making it well-suited for scalable 3D reconstruction in real-world scenarios.
@@
Multi-View Association, Training-Free Visual Perception, Featureless Matching, $\delta$-Overlap Graphs, Graph-Based Object Matching, Epipolar Geometry, 3D Reconstruction
@@ I Introduction
Multi-view multi-object association plays a critical role in 3D reconstruction, aiming to consistently identify and group object instances across multiple camera views. This task is fundamental to a wide range of applications, including autonomous driving [1] and 3D human pose estimation [2, 3, 4]. However, it remains highly challenging due to factors such as occlusions, noisy observations, and imperfect camera calibration, all of which can degrade matching accuracy. Errors in associating 2D observations across views often propagate and significantly impair the quality of the resulting 3D reconstruction. Classical approaches typically rely on distinctive appearance features and geometric constraints such as epipolar consistency, but these methods can struggle in cluttered scenes with numerous objects and limited visual distinguishability.
Recent trends have shifted toward learning-based methods that employ neural networks to associate objects across views [5, 1, 6, 7, 2, 3, 8, 9, 10]. These approaches often integrate appearance cues—such as texture, color, or shape—with spatial and temporal information to resolve correspondence. While effective in visually rich and structured environments, they tend to fail in more difficult cases where appearance cues are ambiguous, degraded, or altogether absent. In particular, scenes containing multiple visually identical instances pose a fundamental challenge: in such settings, even temporal dynamics are insufficient to disambiguate object identities across views.
Figure 1: Demonstration of the C-DOG applied to 130 3D points in a multi-camera system from our benchmark. (a) Multiple visually indistinguishable 3D points are observed within a multi-camera setup. (b) Each camera captures 2D projections of the 3D points. These 2D observations can be obtained using object detectors or keypoint estimators. (c) C-DOG is used to establish correspondences across views by associating multi-view multi-object observations. (d) The resulting correspondences are then used to perform 3D reconstruction.
@@
To overcome this, some approaches dispense with appearance features and instead rely solely on multi-view geometric constraints [4, 11, 12, 13, 14, 15, 16, 17, 18]. Specifically, they leverage known camera poses—which encode the relative position and orientation of each camera—to infer associations between 2D observations across views. These geometric relationships are often encoded as pairwise affinity scores indicating the likelihood of correspondence between detections from different views. A common formulation treats each object detection as a graph node, with weighted edges indicating the strength of geometric consistency between views.
@@
However, these methods are highly sensitive to observation noise and scaled instance counts. As illustrated in Fig. 2, in the simple case of two views and two nearby objects, a noise-free observation yields a distinct association pattern—correct correspondences have zero or near-zero scores—while under noisy conditions, the correct and incorrect associations become nearly indistinguishable. This ambiguity worsens as the number of objects increases. Although graph-based approaches incorporating epipolar constraints or back-projection errors have been proposed, their performance degrades significantly in the presence of noise and as the number of objects increases, limiting their reliability in real-world scenarios.
@@
To address these challenges, we propose a novel training-free multi-view multi-object associator, which integrates connected $\delta$-neighbor-overlap graph (C-DOG) structures with epipolar geometric constraints. This method uses a given threshold $\delta$ to check the likelihood of the correctness of any association pair utilizing the mutual neighbor connections. Designed to be lightweight and easy to integrate, C-DOG acts as a plug-and-play module for bridging object detection or pose estimation with 3D reconstruction, as illustrated in Fig. 1.
@@
Inspired by existing geometry-based multi-view association algorithms, we model each 2D observation from a camera view as an independent node in a graph, and initialize cross-view edges based on epipolar consistency. We then apply a $\delta$-neighbour-overlap clustering  [19] and filtering procedure to construct a graph that captures object correspondences as connected subgraphs, while tolerating missing or noisy edges. This stage filters out erroneous associations based on structural consistency rather than raw epipolar scores, yielding clusters that represent geometrically consistent observations of the same object across views.
@@
To further refine these clusters, we apply a filtering step based on the Interquartile Range (IQR)  [20] of edge consistency scores and the 3D back-projection error, which quantifies the triangulation error of matched points. Our approach is robust to appearance ambiguity and limited view overlap, and significantly increases the feasible number of objects that can be accurately associated under challenging multi-camera conditions.
@@ II Related Works
II-A Deep Learning Multi-view Object Association Methods
@@
Current mainstream approaches to multi-view multi-object association predominantly rely on visual cues, typically employing neural networks to extract implicit features from images [5, 1, 6]. Among them, Transformer-based architectures [21, 22] have demonstrated strong capabilities in aggregating multi-view features by leveraging self-attention mechanisms to capture inter-dependencies within each view’s feature space [1, 7, 2, 3]. To further enhance correspondence matching, spatial camera information—such as epipolar geometry and homography—is often integrated into these pipelines [5, 8, 9]. Temporal cues have also been incorporated to improve consistency across frames. Yu et al.[10] and Zhang et al. [6] exploit features extracted from earlier frames to jointly refine multi-view associations over time. In parallel, Graph Neural Networks (GNNs)[23, 24] have been increasingly adopted for this task, where detections across views are modeled as graph nodes and correspondence estimation is cast as the correspondence group prediction problem[25, 26, 27]. These methods jointly model appearance descriptors and geometric relationships. However, they typically require large-scale, task-specific training data and exhibit reduced robustness in scenarios lacking distinctive visual features—making them less suitable for texture-poor or privacy-sensitive environments. Moreover, deep learning-based methods often incur significant computational overhead. Although some achieve real-time inference in isolation, their integration into larger systems—especially those involving multiple networks or concurrent tasks—can exceed available GPU resources. As such, lightweight, feature-independent alternatives offer practical advantages in both performance and deployability. In addition, reliance on feature-based learning limits adaptability, as these approaches generally require objects with distinguishable visual appearances and demand extensive training, thereby constraining their flexibility in dynamic or unseen environments.
II-B Training-Free Multi-view Object Association Methods
While a handful of deep learning-based methods can handle multi-view multi-object association, a training-free approach would significantly enhance adaptability. Training-free approaches typically treat each observed object as a node and utilize graph methods to establish correspondences. Greedy matchers select the best-scoring edge at each step without backtracking. Classic variants date back to Shafique et al. [28] and Lu et al. [29], while iterative improvements remain popular in real-time pose estimation [30, 31]. These methods typically settle for local optimal solutions and may easily encounter structured errors when the local optimum is incorrect. While these methods aim for efficiency through local optima, other approaches seek global optima. Dong’s method [4] reconstructs 3D points from each candidate pair via triangulation, followed by a voting mechanism to eliminate outliers. Some works formulate the association as a binary pairwise energy problem and resolve it using an ST graph-cut, guaranteeing the global minimum. This idea dates back to Kolmogorov et al.’s multi-camera reconstruction [32] and the volumetric stereo method of Vogiatzis et al. [33], where edge weights are derived from epipolar consistency or 3D back-projection error. Recent systems still incorporate an ST-cut step between upstream detection and downstream optimization, such as in messy tabletop object reconstruction [11] and multi-view multi-person pose reconstruction pipelines [12]. These methods typically require a designated reference view to initiate the matching process. However, they tend to perform poorly in complex scenarios involving numerous objects and are highly sensitive to geometric inaccuracies, such as calibration errors or noisy observations. Similar to ST-cut methods that ensure global consistency, permutation synchronization [34, 13] treats pairwise correspondences as partial permutation matrices and synchronizes them across views. CLEAR [14] adopted this approach in its association pipeline, and Li et al. [15] scaled permutation synchronization to improve computational efficiency. Factorized graph matching (FGM) [16] factorizes the pairwise affinity tensor to avoid costly memory and computational overhead. Composition-based Affinity Optimization (CAO) style extensions [35, 36, 17] further improve both time and space complexity, enabling larger-scale graph matching. Zach proposed a robust bundle adjustment method [37] by lifting the cost function’s dimensionality, allowing soft matching to better handle outliers. RKHS-BA [38] further advanced this by enabling correspondence-free point group association, enhancing pose estimation quality under noisy conditions. Spectral clustering segments an affinity graph using Laplacian eigenvectors [39, 18], but remains sensitive to kernel bandwidth selection and the preset number $k$ of clusters, especially when objects are visually identical. Methods employing connected component clustering [40] construct K-NN graphs and remove edges exceeding a threshold, an approach later adapted for 3D human pose estimation [25, 41].
Although these training-free methods serve as easily integrable modules for multi-view multi-object association, some still rely on visual features represented by feature descriptors such as SIFT [42], resulting in significant degradation in both accuracy and computational efficiency when handling large numbers of points without visual cues. In this work, we propose an efficient yet robust training-free method that utilizes pure spatial geometry (featureless) to establish multi-view multi-object correspondences, thus extending the applicability beyond the limitations of existing approaches.
@@ III-A Epipolar Geometry
where the intrinsic parameter matrices are denoted by $\mathbf{K}_{m}$ and $\mathbf{K}_{m^{\prime}}$, and the extrinsic parameter matrices are denoted by $[\mathbf{R}_{m}\mid\mathbf{T}_{m}]$ and $[\mathbf{R}_{m^{\prime}}\mid\mathbf{T}_{m^{\prime}}]$ respectively. $[\mathbf{T}_{m,m^{\prime}}]_{\times}\in\mathbb{R}^{3\times 3}$ is a skew-symmetric matrix of the cross-product of translation vector $\mathbf{T}_{m,m^{\prime}}$. The $3\times 3$ fundamental matrix has a rank equals two and is skew-symmetric, namely,
@@ III-B Triangulation and Back-Projection Error
Let $\{V_{m};m\in M(\mathbf{r})\}$ be a set of 2D points where $M(\mathbf{r})$ is a subset of indices of camera views such that each 2D point in it is the projection of a 3D point $\mathbf{r}$. Substituting into eq. (6), it can be expressed as
@@
Ideally, one would expect $V_{m}=\hat{V}_{m};m\in M(\mathbf{r})$. However, in practice, their difference
$e_{m}=V_{m}-\hat{V}_{m}$ (8)
is often non-zero. This is known as back projection error (BPE). BPE has a number of potential causes: estimation error of 2D feature points, calibration error of camera poses (intrinsic and extrinsic parameters), and data association error. In this work, we will focus on the data association error due to erroneous association of some 2D observations to a 3D point.
@@ IV Methodology
Figure 4: Overview of C-DOG with an illustrative example with three 3D points and 6 cameras: (a) A multi-camera system observes object points from multiple viewpoints. (b) Required inputs include the camera pose for each view and 2D observed points in pixel coordinates. (c) Initial graph generation: each node represents a 2D point in a specific view, and edges between nodes are weighted by epipolar distances. Only view-wise minimal connections that satisfy the epipolar constraint are retained. (d) Weak connection removal: sparse inter-group connections are identified and eliminated based on node neighborhood overlap density  [19], as they typically represent incorrect associations. (e) Group-level outlier removal: outliers are detected using 3D back-projection error within each group. (f) Final output: associated groups of 2D points across views, where each group corresponds to a distinct 3D object instance.

### v2

# C-DOG: Multi-View Multi-instance Feature Association Using Connected $\delta$-Overlap Graphs thanks: Code available at GitHub: https://github.com/Yunghong/C-DOG

@@ Abstract
Multi-view multi-instance feature association constitutes a crucial step in 3D reconstruction, facilitating the consistent grouping of object instances across various camera perspectives. The presence of multiple identical objects within a scene often leads to ambiguities for appearance-based feature matching algorithms. Our work circumvents this challenge by exclusively employing geometrical constraints, specifically epipolar geometry, for feature association. We introduce C-DOG (Connected $\delta$-Overlap Graph), an algorithm designed for robust geometrical feature association, even in the presence of noisy feature detections. In a C-DOG graph, two nodes representing 2D feature points from distinct views are connected by an edge if they correspond to the same 3D point. Each edge is weighted by its epipolar distance. Ideally, true associations yield a zero distance; however, noisy feature detections can result in non-zero values. To robustly retain edges where the epipolar distance is less than a threshold $\delta$, we employ a Szymkiewicz–Simpson coefficient. This process leads to a $\delta$-neighbor-overlap clustering of 2D nodes. Furthermore, unreliable nodes are pruned from these clusters using an Inter-quartile Range (IQR)-based criterion. Our extensive experiments on synthetic benchmarks demonstrate that C-DOG not only outperforms geometry-based baseline algorithms but also remains remarkably robust under demanding conditions. This includes scenes with high object density, no visual features, and restricted camera overlap, positioning C-DOG as an excellent solution for scalable 3D reconstruction in practical applications.
@@
Multi-View Association, Visual Perception, Featureless Matching, $\delta$-Overlap Graphs, Graph-Based Object Matching, Epipolar Geometry, 3D Reconstruction
@@ I Introduction
Multi-view 3D reconstruction in computer vision  [1, 2] is the process of generating a three-dimensional model of a scene or object from a collection of two-dimensional images captured from different viewpoints. This capability forms the technical foundation for a wide array of applications, including autonomous driving [3], surgery [4], and robotics [5], among others. The classical multi-view 3D reconstruction pipeline typically involves several stages: extraction of feature points from 2D images, associating these 2D feature points from different camera views to corresponding 3D points, constructing an approximated 3D surface based on the estimated 3D point cloud, and finally, reconstructing stereo views by rendering the estimated 3D surface with color/texture information observed from the multi-view images.
Among these steps, feature association [6] presents the most complex challenge, as it seeks to accurately match hundreds or thousands of 2D points to their corresponding 3D locations. Traditionally, this problem is addressed by first matching the visual appearance surrounding each feature point (using feature descriptors) and then imposing geometric constraints, such as homography [7], to simplify computation. It is generally preferable to extract visually distinct features, as this significantly reduces the complexity of feature matching. Recent trends, however, have shifted towards learning-based methods that leverage neural networks to associate objects across views [8, 3, 9, 10, 11, 12, 13, 14, 15]. These approaches often integrate appearance cues—such as texture, color, or shape—with spatial and temporal information to robustly resolve correspondence.
Figure 1: Demonstration of C-DOG applied to 130 3D points within a multi-camera system from our benchmark. (a) Multiple visually indistinguishable 3D points are observed within a multi-camera setup. (b) Each camera captures 2D projections of these 3D points, which can be obtained using object detectors or keypoint estimators. (c) C-DOG is employed to establish correspondences across views by associating multi-view multi-object observations. (d) The resulting correspondences are then utilized to perform 3D reconstruction.
@@
While effective in visually rich and structured environments, existing feature matching approaches often struggle in more challenging scenarios where appearance cues are ambiguous, degraded, or entirely absent. Specifically, scenes containing multiple visually identical instances (multi-instance scenes) pose a fundamental challenge: in such settings, even temporal dynamics may be insufficient to disambiguate object identities across views.
@@
To address these limitations, some approaches abandon appearance features, relying instead solely on multi-view geometric constraints [16, 17, 18, 19, 20, 21, 22, 23, 24]. These methods leverage known camera poses—which encode the relative position and orientation of each camera—to infer associations between 2D observations across views. Such geometric relationships are frequently encoded as pairwise affinity scores, indicating the likelihood of correspondence between detections from different views. A common formulation treats each object detection as a graph node, with weighted edges representing the strength of geometric consistency between views.
@@
However, these geometry-only methods are highly sensitive to observation noise and scale poorly with increasing instance counts. As illustrated in Figure 2, a noise-free observation in a simple case of two views and two nearby objects yields a distinct association pattern where correct correspondences have zero or near-zero scores. In contrast, under noisy conditions, correct and incorrect associations become nearly indistinguishable. This ambiguity intensifies as the number of objects increases. Although graph-based approaches incorporating epipolar constraints or back-projection errors have been proposed, their performance degrades significantly in the presence of noise and a higher number of objects, thus limiting their reliability in real-world applications.
@@
To overcome these challenges, we propose a novel multi-view, multi-instance feature association method that integrates connected $\delta$-overlap graph (C-DOG) structures with epipolar geometric constraints. This method utilizes a predefined threshold $\delta$ to assess the likelihood of correctness for any association pair, leveraging mutual neighbor connections. Designed to be lightweight and easily integrable, C-DOG functions as a plug-and-play module that bridges object detection or pose estimation with 3D reconstruction, as depicted in Figure 1.
@@
Inspired by existing geometry-based multi-view association algorithms, we model each 2D observation from a camera view as an independent node in a graph. Cross-view edges are initially established based on epipolar consistency. We then apply a $\delta$-neighbor-overlap clustering [25] and filtering procedure to construct a graph that captures object correspondences as connected subgraphs, while tolerating missing or noisy edges. This stage filters out erroneous associations based on structural consistency rather than raw epipolar scores, yielding clusters that represent geometrically consistent observations of the same object across views.
@@
To further refine these clusters, we apply an additional filtering step based on the Interquartile Range (IQR) [26] of edge consistency scores and the 3D back-projection error, which quantifies the triangulation error of matched points. Our approach demonstrates robustness to appearance ambiguity and limited view overlap, and significantly increases the feasible number of objects that can be accurately associated under challenging multi-camera conditions.
@@ II Related Works
II-A Multi-view Feature Matching
Feature matching serves as a foundational step for higher-level vision tasks, including localization, mapping, and multi-view multi-object association. Classical approaches, such as ORB [27] and RANSAC [28], offer computational efficiency and robustness in structured environments but often degrade under wide-baseline viewpoints, severe illumination changes, or texture-poor regions. To address these challenges, deep learning–based matchers such as SuperPoint [29], SuperGlue [30], and ClusterGNN [31] leverage learned keypoint detectors, descriptors, and context-aware correspondence reasoning. These methods deliver superior performance in challenging scenarios by jointly optimizing detection and matching but incur substantial computational costs and require extensive training data [6]. Recent advances have moved beyond independent pairwise matching: CoMatcher[32] introduces GNN-based cross-attention and cross-view geometric constraints to collaboratively reason over multiple co-visible images, improving correspondence reliability in wide-baseline and occluded scenes. Similarly, Roessle and Nießner[33] integrate differentiable pose optimization with a GNN-based multi-view matcher, enabling pose errors to guide the learning of correspondences and boosting geometric accuracy. In parallel, optimization-driven alternatives such as MATCHEIG[34] enforce spectral cycle-consistency using eigen-decomposition to refine noisy pairwise correspondences into globally consistent multi-view matches, while UPMGC-SM[35] employs structural graph matching and fusion to align unpaired multi-view data without relying on deep features or paired training data. These developments illustrate a spectrum of approaches, from purely geometric and spectral methods to advanced neural architectures, highlighting the trade-off between scalability, accuracy, and computational demand.
II-B Deep Learning Multi-view Object Association
@@
Building upon feature matching, deep learning–based multi-view multi-object association frameworks integrate visual cues, geometric constraints, and temporal information to produce globally consistent associations across views. Many methods adopt Transformer architectures [36, 37], whose self-attention mechanisms are effective at capturing interdependencies within and across views [3, 10, 11, 12]. To enhance robustness, spatial camera constraints—such as epipolar geometry and homography—are often incorporated to regularize correspondence estimation [8, 13, 14]. Beyond static matching, temporal information has been leveraged to maintain consistency across frames: for example, Yu et al.[15] and Zhang et al.[9] jointly refine multi-view associations by propagating features from preceding frames. Graph Neural Networks (GNNs)[38, 39] are increasingly applied to this problem by modeling detections as graph nodes and formulating correspondence prediction as a structured group prediction task [40, 41, 42]. These methods achieve strong performance in visually complex environments by jointly modeling appearance, geometry, and temporal dependencies. However, their success comes with notable limitations: they require large-scale, task-specific labeled datasets for training, are often computationally expensive, and can exhibit reduced robustness in feature-sparse or privacy-sensitive environments. Moreover, when integrated into larger pipelines involving concurrent tasks, their resource demands can exceed available GPU budgets, limiting their deployability in real-time or resource-constrained settings.
II-C Multi-view Object Association
To overcome the limitations of deep learning–based approaches, training-free methods provide adaptable alternatives that require no task-specific training data. These approaches typically treat each detected object as a graph node and estimate correspondences using spatial or geometric cues. Greedy strategies [43, 44] prioritize local edge scores for fast but suboptimal solutions, while global optimization approaches formulate the problem as an energy minimization task, often solved via graph cuts [45, 46], enabling globally optimal correspondences in settings such as volumetric stereo [17]. Permutation synchronization [47, 19] enforces cross-view consistency by synchronizing partial permutation matrices, with scalable variants such as CLEAR [20] and its successors [21] improving computational efficiency. Factorized Graph Matching (FGM)[22] reduces the memory and computational overhead of multi-view affinity optimization, while Composition-based Affinity Optimization (CAO)[48, 49, 23] refines scalability and time complexity for large-scale problems. Extensions like robust bundle adjustment [50] and RKHS-based optimization [51] introduce soft matching and non-linear lifting to better handle outliers and noisy observations. Spectral clustering approaches [52, 24] segment affinity graphs using Laplacian eigenvectors, while connected component clustering [53] builds K-NN graphs with adaptive pruning for robust segmentation, as later adapted to 3D human pose estimation [40, 54]. Although these approaches may underperform deep learning methods in appearance-rich scenes, they provide lightweight, interpretable, and easily deployable solutions that remain effective in feature-sparse, unstructured, or dynamically changing environments.
@@ III-A Epipolar Geometry
where the intrinsic parameter matrices are denoted by $\mathbf{K}_{m}$ and $\mathbf{K}_{m^{\prime}}$, and the extrinsic parameter matrices are denoted by $[\mathbf{R}_{m}\mid\mathbf{T}_{m}]$ and $[\mathbf{R}_{m^{\prime}}\mid\mathbf{T}_{m^{\prime}}]$ respectively. $[\mathbf{T}_{m,m^{\prime}}]_{\times}\in\mathbb{R}^{3\times 3}$ is a skew-symmetric matrix of the cross-product of translation vector $\mathbf{T}_{m,m^{\prime}}$. The $\times 3$ 3 fundamental matrix has a rank equal to two and is skew-symmetric, namely,
@@
Colinear Case: If the ray from the camera center of the $m^{th}$ camera to a 3D point also intersect another 3D point. We say the camera center and these two 3D points are colinear. All 3D points that are colinear in the $m^{th}$ view shall cast a single 2D point in that camera. Their images cast on another ($m^{\prime th}$) view will all be lying on the same epipolar line associated with that single point in the $m^{th}$ view. Hence, it is possible that for $i^{\prime}_{m^{\prime}}\neq i_{m^{\prime}}$,
$V_{m^{\prime},i^{\prime}_{m^{\prime}}}^{T}\mathbf{F}_{m^{\prime},m}V_{m,i_{m}}=V_{m,i_{m}}^{T}\mathbf{F}_{m,m^{\prime}}V_{m^{\prime},i^{\prime}_{m^{\prime}}}=0$ (5)
@@ III-B Triangulation
Let $\{V_{m};m\in G_{r}\}$ be a set of 2D points where $G_{r}$ is a subset of indices of camera views such that each 2D point in it is the projection of a 3D point $\mathbf{r}$. Substituting into eq. (7), it can be expressed as
@@ III-C Back-Projection Error
Ideally, one would expect $V_{m}=\hat{V}_{m};m\in M(\mathbf{r})$. However, in practice, their difference is often non-zero. One may define a back projection error (BPE) for a specific 3D point $\mathbf{r}$ as:
$BPE(\mathbf{r})=\frac{1}{|G_{r}|}\sum_{m\in G_{r}}\|V_{m,i_{m}}-\hat{V}_{m,i_{m}}\|^{2}$ (9)
BPE may be various due to a number of potential causes: estimation error of 2D feature points, calibration error of camera poses (intrinsic and extrinsic parameters), and feature association error. Nonetheless, it is a metric that rewards consistency between observation (2D feature points) and the model prediction (estimated 3D coordinates view calibrated camera models). Thus, a performance metric to evaluate the quality of multiview feature matching would be the averaged back-projection error
$BPE=\frac{1}{L}\sum_{\mathbf{r}}BPE(\mathbf{r})$ (10)
where $L$ is the total number of detected 3D points whose coordinates are estimated from associated 2D point using triangulation.
@@ IV Methodology
Figure 4: Overview of C-DOG with an illustrative example with three 3D points and 6 cameras: (a) A multi-camera system observes object points from multiple viewpoints. Outliers may exist in the observation. (b) Required inputs include the camera pose for each view and 2D observed points in pixel coordinates. (c) Initial graph generation: each node represents a 2D point in a specific view, and edges between nodes are weighted by epipolar distances. Only view-wise minimal connections that satisfy the epipolar constraint are retained. (d) Weak connection removal: sparse inter-group connections are identified and eliminated based on node neighborhood overlap density  [25], as they typically represent incorrect associations. (e) Group-level outlier removal: outliers are detected using 3D back-projection error within each group. (f) Final output: associated groups of 2D points across views, where each group corresponds to a distinct 3D object instance.